4 A survey shows that 40% of the resident of a particular town read the morning paper, 65% read the afternoon paper, and 10% read both. One resident is randomly selected. What is the probability that this person reads neither the morning nor the afternoon paper? a. b. One resident is randomly selected. What is the probability that this person reads only the morning paper?
Answer #1

This problem can be solved by using the basic lroperprop of set theory like union of sets and intersection of sets.

Selt the total po on so 100 n(A): 40% % 100 40 : Gy the inta unionoe , the numka le n)ho 100-95 00 the gning ro 40-10 ニ30* I have solved the answers in decimal. You can also write them in fractions if the question demands so. Like 0.05= 5/100 = 1/20

**If you found my answer helpful kindly give a positive rating. Thank you.